### v2

# GGBench: A Geometric Generative Reasoning Benchmark for Unified Multimodal Models

Abstract
The advent of Unified Multimodal Models (UMMs) signals a paradigm shift in artificial intelligence, moving from passive perception to active, cross-modal generation. Despite their unprecedented ability to synthesize information, a critical gap persists in evaluation: existing benchmarks primarily assess discriminative understanding or unconstrained image generation separately, failing to measure the integrated cognitive process of generative reasoning. To bridge this gap, we propose that geometric construction provides an ideal testbed as it inherently demands a fusion of language comprehension and precise visual generation. We introduce GGBench, a benchmark designed specifically to evaluate geometric generative reasoning. It provides a comprehensive framework for systematically diagnosing a model’s ability to not only understand and reason but to actively construct a solution, thereby setting a more rigorous standard for the next generation of intelligent systems.
1 Introduction
Artificial intelligence research is undergoing a profound paradigm shift, evolving from the foundational era of passive perception to the more sophisticated stage of contextual understanding [alayrac2022flamingo], and now culminating in the frontier of active generation. This evolutionary leap is epitomized by the recent advent of Unified Multimodal Models (UMMs)[zhang2025unified], such as the groundbreaking GPT-4o[hurst2024gpt] and Nano Banana [google_gemini25_flash_image_2025], which demonstrate an unprecedented ability to fluently process, reason across, and synthesize information from a diverse spectrum of modalities. These UMMs can not only understand complex visual and textual inputs, but also generate rich outputs; notably, advanced systems like Unified-IO 2 [lu2024unified] can even produce images, audio, or actions in response to multimodal instructions. Despite these breakthroughs, existing evaluation benchmarks lag behind – they primarily assess discriminative reasoning (e.g. selecting an answer or classifying a visual input) and thus fail to capture the generative dimension of intelligence [mmeu]. In particular, current tests rarely challenge models to construct solutions (e.g. draw a diagram or formulate a step-by-step proof), which is often essential for domains like geometry. This gap is especially evident in geometric problem-solving, where genuine understanding is inseparable from the ability to plan and generate a diagram through multi-step spatial reasoning [wei2025geoint]. Indeed, current models struggle with this constructive aspect – even advanced multimodal systems have difficulty reliably interpreting complex geometric setups [gllava, trinh2024alphageometry, duan2025codeplot].
The trajectory of evaluation benchmarks over recent years reflects a push toward increasingly complex cognition, yet also reveals a gap in evaluating generative reasoning. Early benchmarks targeted single-modality reasoning or generation, primarily for Large Language Models (LLMs) in purely textual domains. For example, GSM8K [cobbe2021training] and MATH [hendrycks2021measuring] established rigorous tests for mathematical problem-solving in text, requiring models to parse word problems and generate final answers through a step-by-step reasoning chain. Techniques such as chain-of-thought prompting greatly improved performance on these datasets [wei2022chain]. Building on this progress, the community introduced multimodal reasoning challenges that incorporate visual context. Benchmarks like ScienceQA [lu2022learn] require models to combine text understanding with image comprehension [tan2024boosting], and the recent MathVista suite [mathvista] amalgamates 28 math and visual datasets to evaluate mathematical reasoning in visual contexts. These tasks, aimed at Multimodal LLMs (MLLMs), represent a critical step toward contextual understanding – models must interpret a question against an image or figure – but they remain largely discriminative. The model is still to answer or classify based on given content, rather than to create new content. The latest multimodal benchmarks have started to probe both understanding and generation capabilities, but often in a decoupled fashion. Efforts like MME [zhang2021mme], MM-Vet [yu2023mm], and MMBench [liu2024mmbench] evaluate a broad range of perceptual and cognitive tasks to score models on each ability individually. Critically, however, even these comprehensive benchmarks treat understanding and generation as separate modules. This fragmentation means we lack evaluation of the integrated cognitive process – the scenario where a system must simultaneously comprehend, reason, and generate a complex outcome.
Perform generative reasoning demands both cross-modal understanding of abstract concepts and the structured, stepwise generation of a coherent solution. We propose that geometric construction problems provide an ideal testbed for evaluating such integrated understanding–generation capabilities within a unified setting. Such problems inherently require a model to:
Parse and comprehend abstract natural language instructions with domain-specific constraints;
Formulate a multi-step plan grounded in formal geometric principles;
Generating precise figures that satisfies the given constraints.
Performance on geometric construction tasks is directly verifiable, offering a clear and objective criterion for evaluation. Success in such tasks engages the full spectrum of intelligence: language understanding, mathematical reasoning, long-horizon planning, and visual generation. To enable this rigorous evaluation, we introduce GGBench, a comprehensive benchmark where each problem instance consists of precisely aligned text, code, and image modalities, as shown in Figure 2, facilitating a systematic analysis of the entire multimodal generative reasoning process.
2 Related Work
2.1 The Evolution of Mathematical Reasoning Benchmarks
The evaluation of mathematical reasoning in AI has progressed from discriminative question answering toward more generative, process-oriented tasks. Early benchmarks such as GeoQA [chen2021geoqa] provides valuable testbeds but narrowly targets plane figures and evaluates models on answer accuracy. Several recent benchmarks like MATH-Vision (MATH-V) [mathv], MathVista [mathvista], and MathVerse [mathverse] have broadened the scope of evaluation by combining visual and textual information. Specialized datasets like PolyMath [polymath] evaluates general cognitive reasoning in multimodal settings, requiring models to combine visual pattern recognition with mathematical logic. MathScape [mathscape] uses real-world photographs for math questions. GeoEval [zhang2024geoeval] compiled a comprehensive suite of geometry question. SolidGeo [solidgeo] focuses on solid/3D geometry, whereas VisAidMath [visaidmath] evaluates the use of visual aids such as drawing auxiliary lines in geometry. In a different vein, MMSciBench [ye2025mmscibench] evaluates multimodal scientific problems, covering both math and physics questions in textual and text+image formats. Beyond static images, researchers have explored temporal problem-solving paradigms: VideoMathQA [videomathqa] assesses long-horizon mathematical reasoning in educational videos, and NewtonBench [newtonbench] challenges LLM-based agents to discover scientific laws through interactive experimentation. MathOPEval [mathopeval] requires models to perform fine-grained visual operations by generating executable code like constructing, modifying, or annotating a diagram. DynaMath [dynamath] introduces dynamic variations of visual math problems to test robustness. MaRVL-QA [marvlqa] uses unlabeled function plots and geometry transformations as inputs that minimizes textual cues.
Crucially, as researchers identified the limitations of purely result-oriented evaluation, new benchmarks have placed increasing emphasis on the reasoning process itself. MM-MATH [sun2024mm] exemplifies this shift by augmenting outcome evaluation with process evaluation. We-Math [wemath] draws inspiration from human problem-solving principles and decomposes problem into sub-problems aligned with knowledge concepts. Math2Visual [wang2025generating] introduces the task of generating pedagogically useful diagrams for math word problems. A model is given a math problem in text and must produce a diagram or visual illustration that meaningfully represents the problem’s content. Most recently, Geoint-R1 [wei2025geoint] integrates auxiliary element construction with formal verification, and MathCanvas [shi2025mathcanvas] enabled a unified multimodal model [bagel] to learn ”when and how” to draw as part of its reasoning process. Despite this rich and rapidly evolving landscape, a critical gap persists. The majority of existing evaluations, even those that are process-oriented, ultimately reduce to either checking a final numerical or categorical answer or scoring a model’s generated text against a reference solution. A fundamental challenge remains unaddressed: the lack of a benchmark that holistically evaluates a model’s ability to generate a complete, multi-step, and formally verifiable geometric construction from scratch.
2.2 The Rise of Unified Multimodal Models
UMMs signifies a pivotal architectural shift, not merely in the unification of modalities, but more fundamentally in the unification of understanding and generation capabilities within a single framework [zhang2025unified]. This paradigm moves beyond specialized models towards a generalist capable of both interpreting multimodal inputs and generating structured outputs. Recent examples of UMMs include GPT-4o [hurst2024gpt] and Gemini 2.5 [comanici2025gemini], as well as powerful open-source alternatives such as Janus series [janus, chen2025janus, ma2025janusflow], Qwen-VL [bai2023qwen, wang2024qwen2], and OmniBridge [omnibridge], are all designed to seamlessly transition between comprehension and synthesis. Other efforts include Apple’s MM1 [mckinzie2024mm1] and Google’s Nano Banana (Gemini 2.5 Flash Image) [google_gemini25_flash_image_2025]. There are also hybrid systems like MM-ReAct [yang2023mm], which augment an LLM with external visual tool interfaces rather than a single unified network. Together, these developments mark a shift from modality-specific AI toward end-to-end multimodal intelligence, with UMMs learning unified representations that support cross-modal understanding and generation.
On the one hand, the understanding capabilities of these models have been rigorously validated across a vast landscape of benchmarks. They have achieved exceptional performance on testbeds like MMMU [yue2024mmmu] for expert-level knowledge, MME-Unify [xie2025mme] for holistic perception and cognition, and MathVista [mathvista] for visual mathematical reasoning. Models explicitly designed for deep comprehension, such as Query-Kontext [querykontext] and MAGUS [magus], further demonstrate this proficiency. On the other hand, the generative prowess of UMMs is rapidly advancing, moving far beyond simple image captioning. Models like Bagel [bagel] and its successor Hyper-Bagel [hyperbagel] showcase sophisticated generative abilities honed by novel training techniques. Application-specific models like ChartSketcher [chartsketcher] can synthesize data visualizations from text, while generative models like Lingshu [lingshu], VARGPT [vargpt], and HaploOmni [haploomni] exhibit advanced faculties for creating rich, context-aware multimodal content. The development of sophisticated evaluation frameworks like UniEval [li2025unieval] and reward models such as Skywork-VL Reward [skyworkvlreward] reflects a growing effort to quantify and guide complex generative behaviors. Despite these parallel advancements in understanding and generation, a critical gap persists in existing evaluation. Benchmarks have largely focused on assessing these two functions as separate capabilities. Consequently, there remains a lack of a framework for evaluating a model’s ability to integrate these functions in a generative reasoning task—that is, tasks that require a model to first deeply comprehend multimodal premises and then generate a structured, logically sound, and verifiable artifact as a solution.
2.3 Code-based Evaluation for Reasoning Tasks
To enable rigorous evaluation, a growing body of work adopts code as a structured mechanism for assessment, leveraging the verifiable nature of code over free-form natural language explanations. MathCoder-VL [mathcodervl] employs image-to-code supervision to ensure precise visual–textual alignment, while MATP-BENCH [matpbench] integrates formal theorem proving in Lean, Coq, and Isabelle for verifiable multimodal logic. VeriEquivBench [veriequivbench] proposes an equivalence-based validation metric for verifying consistency between generated code and specifications. InternLM-Math [internlmmath] unifies solving and verifying under a single framework with interleaved reasoning and code execution. DeepMath-103K [deepmath103k] provides a large, contamination-free dataset with verifiable answers for reinforcement learning with rule-based rewards. MathQ-Verify [mathqverify] targets problem validity itself through a structured verification pipeline. CMMaTH [cmmath] introduces a large Chinese multimodal math benchmark and an open-source automatic grader, while MARIO Eval [marioeval] standardizes automated evaluation across mathematical datasets. U-MATH [umath] assess university-level reasoning and meta-evaluate LLM judges. In parallel, QuesCo [quesco] captures holistic mathematical intent through contrastive pretraining.
While recent multimodal datasets [mathvista, wemath, mathscape, mathv, mathllava, mvmath] have advanced the field, they seldom enforce a verifiable alignment between the natural language reasoning steps. To bridge this gap, GGBench extends the code-based evaluation to the visual geometric domain, enabling a holistic verification pipeline that tightly aligns textual reasoning, code execution, and visual output. This multimodal, verifiable setup allows GGBench to move beyond pattern matching and surface-level correctness.
3 The GGBench Benchmark
3.1 Overview
We introduce GGBench, a geometric generative reasoning benchmark designed to evaluate UMMs through end-to-end construction tasks. Unlike conventional discriminative benchmarks that evaluate models solely on final answer selection or numerical computation, GGBench operationalizes the assessment of whether a system can genuinely construct a geometrically valid solution.
A defining feature of GGBench is its tri-modal data structure, where each sample comprises precisely aligned text, code, and image components. This multi-faceted design enables a uniquely versatile and diagnostic evaluation framework. For instance, the complete, end-to-end task of generating a final image from a textual prompt serves as the primary test for the integrated generative reasoning of UMMs. Concurrently, the benchmark can be decoupled to probe specific capabilities: the natural language reasoning steps (text) and their translation into executable programs (code) can be used to assess the logical planning and code generation faculties of LLMs. Furthermore, the provided GeoGebra code acts as an unambiguous, machine-verifiable ground truth, offering a deterministic method for verifying the geometric correctness and precision of any generated figure. An overview of the multi-modal dataset construction pipeline is shown in Figure 3, which defines the stage annotations (a)–(f) referenced throughout the paper.
3.2 Dataset construction
Data collection and candidate pooling We manually search the web for public geometry problems and assemble a pool that spans classical constructions and contest-style tasks (Figure 3, stage (a)). To scale selection beyond purely manual effort, we use an LLM for assisted tagging and filtering (LLM-i in the figure) and then conduct human checking to retain items that are suitable for construction—i.e., problems with unambiguous geometric dependencies and diagrammatically actionable conditions (Figure 3, stage (b)). This stage emphasizes diversity in objects and relations (e.g., circles, tangents, incenters, parallels) while discarding under-specified prompts.
Prompt design and problem adaptation To elicit end-to-end construction in GeoGebra only, we design a composite prompt that pairs a textual specification with exemplar GGB code (Figure 3, stage (c)). In parallel, we prepare a reformulation prompt family that converts the selected problems into construction-oriented statements with explicit auxiliary objects and ordered dependencies (Figure 3, stage (d)). In practice, GPT-5 [OpenAI2025GPT5SystemCard] serves as the primary authoring model during adaptation (LLM-ii), while GGB-oriented prompt templates are informed by prior engineering with Gemini 2.5 Pro [comanici2025gemini] to improve syntactic fidelity. The outcome is a set of adapted problems whose reasoning steps align one-to-one with executable GGB operations.
Solution generation and visualization Given an adapted problem and its data-generation prompts, GPT-5 [OpenAI2025GPT5SystemCard] (LLM-iii) produces synchronized outputs: (i) stepwise reasoning text, (ii) executable GGB code, and (iii) a rendered diagram obtained by running the code (Figure 3, stage (e)). This stage focuses on faithful construction: steps materialize auxiliary objects (e.g., angle bisectors, tangents, parallels), code reflects those steps, and the rendered figure visualizes the target configuration. At this point we obtain roughly draft instances prior to filtering.
Automated screening and expert finalization The final stage involved a rigorous two-tier filtering process to ensure dataset quality. First, we perform LLM-based quality control (LLM-iv) along three axes—code executability, logical consistency of the construction process, and diagram correctness with respect to the intended figure—to remove low-quality drafts (Figure 3, stage (f)). Subsequently, all surviving instances underwent a final review by domain experts who verified their geometric correctness, the sufficiency of the construction, and the precise consistency across the text, code, and image modalities. After this two-tier filtering, GGBench retains high-quality items with tightly aligned text–code–image triplets. A detailed statistical analysis of the dataset is provided as below.
| Dataset size | |
| Total problems | 1,411 |
| Problem types | |
| Straightedge-and-compass construction | 798 |
| Geometric transformation construction | 426 |
| Analytic construction | 187 |
| Difficulty levels | |
| Easy | 298 |
| Medium | 816 |
| Hard | 297 |
| Images per problem | |
| 3 images | 7 |
| 4 images | 227 |
| 5 images | 860 |
| 6 images | 283 |
| 7 images | 34 |
| Total images | 7,165 |
| Question length (tokens) | |
| Minimum | 68 |
| Maximum | 483 |
| Average | 189.83 |
3.3 Dataset Analysis
Dataset Composition and Scale
GGBench comprises 1,411 GeoGebra-based construction problems that integrate reasoning, code, and visual outputs. Table 1 summarizes overall dataset statistics, including problem types, difficulty levels, image counts, and question lengths. Most problems belong to the medium difficulty (57.83%), balancing solvability and reasoning complexity. Straightedge-and-compass constructions dominate (56.6%), followed by geometric transformation and analytic constructions. Each problem includes between 3 and 7 diagrams, averaging 5.08 images per problem, providing dense visual evidence for evaluating generative reasoning.
| Category | Count |
| Basic Constructions | 1,063 |
| Circle Properties & Constructions | 931 |
| Geometric Transformations | 376 |
| Triangle Properties & Constructions | 280 |
| Applications of Geometric Theorems | 218 |
| Polygon Properties & Constructions | 107 |
| Measurement & Ratios | 92 |
| Locus Constructions | 30 |
| Total category tags | 3,097 |
Category Distribution and Difficulty
GGBench covers eight major categories of geometric reasoning, as detailed in Table 2. Each problem may involve multiple skills, resulting in 3,097 category annotations across the dataset. As shown in Figure 4, Basic Constructions and Circle Properties & Constructions are predominant across all difficulty levels, while advanced reasoning types—such as Applications of Geometric Theorems and Measurement & Ratios—become more frequent in the hard subset. This stratification ensures balanced coverage from elementary to advanced geometric reasoning.
Category Composition and Cognitive Coverage
The distribution in Table 2 reflects a deliberate cognitive gradient. The majority of tasks emphasize procedural reasoning (e.g., bisectors, parallels, and tangents), providing a stable foundation for evaluating geometric comprehension. Intermediate categories, including Transformations and Triangles, introduce abstraction through relational constraints and inter-object dependencies. Higher-order categories—such as Theorem Applications and Measurement & Ratios—demand symbolic reasoning, multi-step dependencies, and precision alignment between text and geometry. This stratification enables a fine-grained evaluation of a model’s capabilities, from foundational geometric execution to advanced, abstract problem-solving.
Token Length and Reasoning Path
We report question length in tokens. As shown in Table 1, question length ranges from 68 to 483 tokens, averaging 189.83 tokens. This range reflects the multi-step reasoning nature of GGBench tasks, requiring the model to plan, describe, and validate geometric constructions. On average, each problem contains 5.08 rendered images (7,165/1,411), and carries 2.20 category tags (3,097/1,411), indicating that GGBench is both visually dense and multi-skill by design. The five-image setting is the most common (60.9%; 860/1,411), providing ample visual evidence for stepwise construction. While Basic Constructions and Circle Properties & Constructions dominate across the corpus, Figure 4 further shows a progressive shift toward higher-order skills (e.g., Applications of Geometric Theorems, Measurement & Ratios) in the hard subset.
3.4 Comparison with Existing Benchmarks
Table 3 situates GGBench among representative multimodal math benchmarks along two groups of axes: (i) capability coverage—the proportion of samples that exercise understanding, generation, and multi-step reasoning; and (ii) modality supervision—whether problems provide aligned text, images, and code. We normalize modality coverage as a binary indicator in the table (✓/✗) for readability.
| Benchmark | Understanding | Generation | Multi-step | Text | Image | Code |
| MathVista [mathvista] | 49.9 | 34.8 | 21.5 | ✓ | ✓ | ✗ |
| MathVerse [mathverse] | 61.3 | 52.7 | 45.2 | ✓ | ✓ | ✗ |
| MM-MATH [sun2024mm] | 31.0 | 26.5 | 82.0 | ✓ | ✓ | ✗ |
| MathScape [mathscape] | 44.2 | 39.8 | 18.5 | ✓ | ✓ | ✗ |
| GeoEval [zhang2024geoeval] | 46.0 | 25.0 | 33.3 | ✓ | ✓ | ✗ |
| WE-MATH [wemath] | 58.7 | 51.9 | 42.6 | ✓ | ✓ | ✗ |
| MMSciBench [ye2025mmscibench] | 55.4 | 32.8 | 40.2 | ✓ | ✓ | ✗ |
| MATH2VISUAL [wang2025generating] | 48.0 | 37.0 | ✗ | ✓ | ✓ | ✗ |
| PolyMath [polymath] | 54.6 | ✗ | ✗ | ✓ | ✗ | ✗ |
| SOLIDGEO [solidgeo] | 49.5 | 18.7 | 6.7 | ✓ | ✓ | ✗ |
| GGBench | 100.0 | 100.0 | 100.0 | ✓ | ✓ | ✓ |
Generative construction vs. answer selection
Most benchmarks [mathvista, mathscape] emphasize recognition or answer selection. Even when generation is present [mathverse], outputs are rarely required to be constructive artifacts that satisfy constraints. By contrast, every GGBench instance requires producing diagrams with formal constraints, shifting evaluation from choosing an answer to constructing evidence.
Tri-modal alignment and verifiability
A central gap in prior work is the absence of code supervision. It is difficult to verify whether a model’s reasoning aligns with the final diagram. GGBench provides 100% alignment between text (stepwise plan), code (executable construction), and image (rendered diagram), enabling process-level judging, program executability checks, and geometry-aware verification within a single benchmark.
Multi-step supervision
While some datasets include multi-step elements [sun2024mm, mathverse], they do not couple each step with executable constructions, limiting verifiability. GGBench enforces multi-step reasoning at 100% and ties steps to code, ensuring that auxiliary objects and dependencies are operationalized rather than narrated.
Modal breadth and heterogeneity.
Most benchmarks offer full text and image coverage (100%/100%), with exceptions such as MMSciBench [ye2025mmscibench] (90%/40%) and PolyMath [polymath] (100%/0%). GGBench maintains complete modality support (text/image/code all 100%), so reasoning, rendering, and execution can be evaluated cohesively.
Together, these dimensions position GGBench as a construction-centric benchmark that unifies understanding, reasoning, and generation under a verifiable framework. It complements prior understanding-oriented datasets, offering a rigorous testbed for the next generation of unified multimodal reasoning models.
4 Experiment
Setup.
We evaluate existing Unified Multimodal Models on GGBench to assess their generative geometric reasoning capabilities. Each model receives the problem text and reference image as input. To ensure deterministic outputs, we fix the temperature at 0.0 during inference. This eliminates stochasticity and enforces consistent step-by-step reasoning. All code generation adheres strictly to the GeoGebra command syntax to ensure executability and structural correctness. Inference is parallelized for efficiency. All models are evaluated under identical settings with unified prompts and processing pipelines. Full prompt templates are detailed in Appendix.
4.1 Baselines
We evaluate models under two complementary tracks that correspond to distinct architectural paradigms: (A) end-to-end UMMs that directly generate diagram images from natural-language prompts, and (B) LLMs/LRMs that first produce explicit, executable construction code which is then rendered into diagrams. This dual-track evaluation allows us to quantify and analyze the gap between UMMs’ immediate visual generation and the geometrically grounded, code-driven constructions produced by LLMs/LRMs.
Track A: End-to-end UMMs These systems take natural-language specifications and output diagram images step-by-step. We include: Qwen-Image [wu2025qwen], Seedream 4.0 [seedream2025seedream], Janus [janus], BAGEL [bagel], and Nano Banana [google_gemini25_flash_image_2025]. Among them, Qwen-Image and Seedream 4.0 produce only final images without stepwise explanations, whereas the other models generate multi-stage visual outputs for step evaluation.
Track B: Planning Code Render These models first produce a textual plan and then emit GeoGebra code to render into images: GPT-4o [hurst2024gpt], GLM-4.5V [hong2025glm], Qwen3VL-235B-A22B [yang2025qwen3], GPT-4 [achiam2023gpt], GPT-5 [OpenAI2025GPT5SystemCard], Claude Sonnet 4.5 [Anthropic2025ClaudeSonnet4_5], Gemini-2.5-Pro [comanici2025gemini], DeepSeek-R1 [liu2024deepseek], DeepSeek-V3.1 [liu2024deepseek], and Qwen3-14B [yang2025qwen3].
4.2 Metrics
To evaluate geometric generative reasoning, We adopt a four-stage protocol: (1) Planning, (2) Middle Process, (3) Final Result, and (4) Overall Scores. All automatic scores are produced by a VLM GPT-4o using fixed prompts.
(1) Planning (VLM-T) measures the model’s ability to formulate a coherent step-by-step plan in natural language before any code or drawing is produced. Given the textual reasoning output, the VLM judge scores each response along three dimensions: (i) Logical coherence, (ii) Step completeness, and (iii) Geometric correctness. Each criterion is rated on a 1–5 scale and rescaled to [0,100]. Details of the judging rubric and prompt design appear in Appendix 8.2.
(2) Middle process (VLM-I-Mid) To evaluate intermediate reasoning consistency, we concatenate all intermediate construction images into a chronological panel and feed it to the same VLM judge. The model assesses: (i) Step accuracy and (ii) Process consistency. Implementation and prompt details are provided in Appendix 8.4.
(3) Final result (VLM-I-Res) assesses the geometric correctness of the final diagram relative to the reference solution. In addition, we also report pixel-level metrics. Details are in Appendix 8.3.
(4) Overall scores The overall VLM score (VLM–I) is computed as the mean of the intermediate VLM-I-Mid and final scores VLM-I-Res. Human raters follow the same rubric as the VLM judge; inter-rater consistency and calibration procedures are detailed in Appendix 8.6. We observe a strong Pearson correlation () between VLM and human scores, validating the high reliability of our automated evaluation for this task.
4.3 Main Results
Table 4.3 summarizes the overall performance across both tracks. End-to-end UMMs remain significantly behind code-driven models in nearly all evaluation dimensions. Even the strongest UMM, Nano Banana, ranks only in the middle tier of code-based systems, illustrating that direct visual generation still struggles to enforce geometric constraints. C”ode-level metrics of LLMs/LRMs are reported in Table 5.
The planning stage (VLM-T) evaluates a model’s capacity to reason step-by-step before any code or drawing is produced. Interestingly, Nano Banana achieves planning scores comparable to several LLMs/LRMs. However, models such as GPT-4o and GLM-4.5V, despite generating coherent plans, often fail to produce executable code, limiting their downstream visual accuracy. This gap demonstrates that high-level reasoning alone is insufficient without grounding in executable geometry.
Intermediate and final constructions (VLM–I) Across both the intermediate (VLM–Imid) and final (VLM–Ires) stages, code-driven models consistently outperform end-to-end generators. Explicitly reasoning through textual plans and producing executable constructions leads to markedly higher geometric correctness and visual coherence. While robust planning does not always guarantee flawless drawings, weak or inconsistent reasoning almost always results in invalid geometry—reinforcing the necessity of tightly coupled reasoning–generation pipelines. Moreover, purely pixel-based metrics (PSNR, SSIM, LPIPS) show only weak correlation with geometric validity: high perceptual similarity can mask structural errors such as misaligned intersections or missing tangencies. Hence, evaluation based on verifiable geometry, rather than appearance alone, offers a more reliable measure of multimodal reasoning quality.
| Model | Planning | Middle Process | Final Result | Overall Scores | ||||
| VLM-T ↑ | VLM-I-Mid ↑ | VLM-I-Res ↑ | LPIPS ↓ | PSNR ↑ | SSIM ↑ | VLM-I ↑ | Human ↑ | |
[…]
| Qwen-Image [wu2025qwen] | - | - | 22.75 | 56.39 | 58.23 | 48.06 | 22.75 | 25.56 |
| Seedream 4.0 [seedream2025seedream] | - | - | 24.45 | 51.06 | 59.44 | 56.44 | 24.45 | 37.56 |
| Janus [janus] | 33.85 | 21.69 | 19.76 | 57.74 | 57.76 | 60.97 | 20.73 | 19.46 |
| BAGEL [bagel] | 23.07 | 21.84 | 19.99 | 57.07 | 61.78 | 58.82 | 20.91 | 20.12 |
| Nano Banana [comanici2025gemini] | 58.54 | 44.83 | 22.81 | 51.85 | 64.53 | 59.51 | 33.82 | 45.75 |
[…]
| GPT-4o [hurst2024gpt] | 59.73 | 26.19 | 2.66 | 95.43 | 5.45 | 5.69 | 14.43 | 23.04 |
| GLM-4.5V [hong2025glm] | 53.32 | 25.63 | 5.02 | 52.91 | 12.19 | 12.94 | 15.33 | 30.14 |
| Qwen3-14B [yang2025qwen3] | 58.65 | 39.30 | 12.97 | 78.81 | 23.92 | 24.81 | 26.13 | 38.23 |
| Gemini 2.5 Pro [comanici2025gemini] | 38.50 | 37.41 | 15.80 | 68.39 | 37.17 | 39.73 | 26.61 | 44.68 |
| DeepSeek-R1 [liu2024deepseek] | 61.16 | 62.42 | 20.48 | 66.06 | 37.94 | 37.59 | 41.45 | 49.55 |
| GPT-4 [achiam2023gpt] | 55.66 | 50.99 | 15.10 | 67.35 | 35.26 | 38.31 | 33.04 | 55.99 |
| Qwen3-VL [yang2025qwen3] | 56.40 | 49.55 | 23.94 | 39.40 | 52.33 | 58.71 | 36.74 | 66.77 |
| DeepSeek-V3.1 [liu2024deepseek] | 60.24 | 73.13 | 26.41 | 57.21 | 48.33 | 50.12 | 49.77 | 68.12 |
| Claude Sonnet 4.5 [Anthropic2025ClaudeSonnet4_5] | 61.19 | 77.92 | 30.29 | 52.22 | 51.74 | 50.52 | 54.11 | 72.12 |
| GPT-5 [OpenAI2025GPT5SystemCard] | 62.01 | 76.79 | 37.36 | 49.65 | 54.80 | 59.49 | 57.08 | 83.06 |
| Model | Pass@1 ↑ | BLEU Mean ↑ | RUBY Mean ↑ | ROUGE-L Mean ↑ | chrF Mean ↑ | EditDist Mean ↓ |
| GPT-4o [hurst2024gpt] | 7.87 | 8.33 | 12.85 | 27.41 | 29.17 | 78.92 |
| GLM-4.5V [hong2025glm] | 14.25 | 10.28 | 17.53 | 30.48 | 38.32 | 79.61 |
| Qwen3-14B [yang2025qwen3] | 34.30 | 14.04 | 23.85 | 38.07 | 44.36 | 76.89 |
| Gemini 2.5 Pro [comanici2025gemini] | 32.67 | 24.21 | 40.56 | 50.50 | 65.87 | 71.14 |
| DeepSeek-R1 [liu2024deepseek] | 54.36 | 18.46 | 33.68 | 42.69 | 60.11 | 73.64 |
| GPT-4 [achiam2023gpt] | 50.53 | 17.31 | 28.54 | 44.62 | 51.25 | 70.34 |
| Qwen3-VL- [yang2025qwen3] | 69.31 | 23.24 | 38.61 | 50.73 | 63.51 | 70.24 |
| DeepSeekV3.1 [liu2024deepseek] | 70.09 | 20.44 | 35.14 | 47.36 | 59.24 | 70.87 |
| Claude Sonnet 4.5 [Anthropic2025ClaudeSonnet4_5] | 75.34 | 22.04 | 36.77 | 49.68 | 60.28 | 70.81 |
| GPT-5 [OpenAI2025GPT5SystemCard] | 79.02 | 18.87 | 39.92 | 44.08 | 62.44 | 74.13 |
Code-based analysis (Table 5) At the program level, GPT-5 achieves the highest execution accuracy (pass@1=79.02), followed closely by Claude Sonnet 4.5 (75.34) and DeepSeek-V3.1 (70.09). Interestingly, code-similarity metrics (e.g., BLEU or ROUGE-L) do not necessarily predict executable success: Gemini-2.5-Pro attains high textual similarity yet a lower Pass@1, confirming that surface-level resemblance does not imply geometric equivalence. This observation validates the necessity of execution-based evaluation in GGBench.
Summary of findings In summary, reasoning-grounded models that produce and execute code achieve substantially higher geometric correctness, interpretability, and human-rated quality. GPT-5 attains the best overall performance (VLM–I=57.08, Human=83.06), followed by Claude Sonnet 4.5 and DeepSeek-V3.1. Despite recent advances, purely generative UMMs remain limited in enforcing spatial constraints, emphasizing the importance of explicit, verifiable construction pipelines for next-generation multimodal reasoning systems.
4.4 Performance by Construction Category
Figure 5 reports the average VLM–I scores across eight construction categories in GGBench, revealing distinct performance patterns among models. GPT-5 [OpenAI2025GPT5SystemCard] achieves the highest overall accuracy, consistently outperforming other models across all categories. Its strength lies in maintaining both symbolic precision (accurate translation from plan to executable code) and spatial coordination (faithful geometric realization). Claude-Sonnet-4.5 [Anthropic2025ClaudeSonnet4_5] ranks second, closely followed by DeepSeek-V3.1 [liu2024deepseek], particularly excelling in tasks requiring geometric regularity such as Geometric Transformations and Circle Properties & Constructions. These models exhibit stable reasoning across categories, indicating robust integration between linguistic and spatial modalities. Qwen3-VL [yang2025qwen3] demonstrates strong performance on structurally constrained categories—such as Basic Constructions and Triangle Properties & Constructions—where relationships among objects are more rule-based, but its scores decline in Measurement & Ratios and Applications of Theorems, which require quantitative and theorem-level reasoning. This contrast suggests that while Qwen3-VL effectively grounds syntactic geometry, it struggles with deeper symbolic arithmetic integration. At the lower end, vision-first models such as GPT-4o [hurst2024gpt], GLM-4.5V [hong2025glm], and Qwen-Image [wu2025qwen] consistently underperform, with mean scores below 23 across most categories. Their weakness is especially evident in tasks requiring explicit relational constraints or quantitative precision, implying insufficient geometric grounding despite competent visual synthesis. Across all models, the most challenging categories are Measurement & Ratios and Applications of Theorems, where performance drops by 10–15 points relative to simpler procedural tasks. These categories require not only visual generation but also algebraic and deductive reasoning, emphasizing the need for symbolic–geometric alignment. Conversely, Basic Constructions and Circle Properties yield the highest scores overall, indicating that most models can handle canonical constructions but fail to generalize to theorem-driven or quantitative reasoning.
Discussion
The category-wise trends highlight the diagnostic value of GGBench. By decomposing performance along interpretable geometric skills, the benchmark exposes which reasoning components—procedural, transformational, or quantitative—limit current multimodal systems. The results confirm that achieving high-fidelity diagram generation alone does not guarantee correct geometric reasoning, underscoring the necessity of construction-level, verifiable evaluation frameworks.
4.5 Performance by Question Type
Figure 6 presents the performance of all evaluated models across three geometric task types: analytic construction (AC), geometric transformation construction (GTC), and straightedge-and-compass construction (SCC), measured using the unified VLM-I metric. Clear distinctions emerge across categories. Models generally perform best on SCC tasks that emphasize rule-based geometric procedures, where GPT-5 [OpenAI2025GPT5SystemCard] reaches 72.54, followed closely by Claude-Sonnet-4.5 [Anthropic2025ClaudeSonnet4_5] and DeepSeek V3.1 [liu2024deepseek], both above 66. Performance on GTC tasks is moderately lower: Claude-Sonnet-4.5 records 66.26 and DeepSeek V3.1 achieves 53.53, reflecting the increased complexity of maintaining geometric invariants under rotation or reflection. Analytic construction tasks yield the lowest results, with GPT-5 scoring 51.90 and DeepSeek V3.1 reaching 46.07, likely due to the greater spatial flexibility and reduced structural constraints of such problems. UMMs such as Janus [janus], BAGEL [bagel], and Seedream [seedream2025seedream] remain consistently weaker due to the absence of symbolic reasoning, while text-only systems like DeepSeek V3.1 show competitive performance on structure-driven categories through accurate code synthesis. Overall, these findings indicate that model effectiveness is shaped jointly by modality alignment and the specific geometric reasoning paradigm required by the task.
Discussion
Overall, performance varies systematically with the degree of symbolic structure embedded in the task. Tasks governed by explicit geometric rules (SCC) favor models with strong language-to-code translation ability, whereas transformation and analytic tasks expose weaknesses in maintaining geometric invariants and algebraic reasoning. These results underscore that multimodal effectiveness in GGBench is shaped jointly by a model’s modality alignment and by the specific reasoning paradigm—procedural, transformational, or analytic—required by the construction type.
4.6 Performance by Difficulty Level
Figure 7 presents the VLM-I scores across three difficulty levels—Easy, Medium, and Hard—providing a comprehensive view of how reasoning complexity impacts model performance. The majority of systems perform best on Easy problems and experience a steady decline as task difficulty increases, confirming the progressive design of GGBench. This pattern is especially pronounced for models like Qwen3-14B [yang2025qwen3] and GPT-4 [achiam2023gpt], which drop more than 20 points from Easy to Hard, suggesting that their performance is disproportionately sensitive to longer reasoning chains and higher geometric abstraction. In contrast, GPT-5 [OpenAI2025GPT5SystemCard] maintains strong results across all difficulty tiers, scoring 72.70 on Easy and 66.77 on Hard, highlighting its robustness in both planning and execution. Claude-Sonnet-4.5 [Anthropic2025ClaudeSonnet4_5] follows a similar trend, while DeepSeek V3.1 [liu2024deepseek] notably surpasses both on Easy problems with a peak of 73.57. UMMs like Nano Banana [comanici2025gemini], Janus [janus], and Seedream [seedream2025seedream] show more stable but consistently lower performance across difficulties, indicating weaker adaptation to reasoning complexity despite visual fluency. These observations reinforce that difficulty scaling in GGBench effectively distinguishes between symbolic robustness and shallow pattern matching, making VLM-I a reliable diagnostic signal for geometric reasoning across complexity levels.
5 Error Analysis
To better understand failure patterns in geometric reasoning, we analyze incorrect outputs from benchmark evaluation and identify four common error types: (i) geometric logic errors, due to misapplied theorems; (ii) structural and contextual errors, involving confusion in figure containment or spatial relationships; (iii) conflation of construction and numeric goals; and (iv) code-level failures from syntax misuse or reserved keyword conflicts. Figure 8 presents a structural and contextual error. The task asks for a rectangle inscribed in a circle, but the model produces a circle inside a square, reversing the intended containment. This highlights a failure to ground symbolic reasoning in geometric constraints, where surface-level textual correctness leads to invalid constructions.
6 Conclusion
We introduce GGBench, the first benchmark explicitly designed to evaluate geometric generative reasoning—the ability of a model to not only understand and reason but also to construct verifiable solutions. Each problem in GGBench integrates natural-language instructions, executable GeoGebra code, and rendered diagrams, forming a tri-modal, interpretable, and fully verifiable testbed. Our unified evaluation protocol jointly measures textual planning quality, code executability and geometric equivalence, and visual diagram fidelity, providing a holistic view of multimodal reasoning competence. Comprehensive experiments across state-of-the-art UMMs and LLMs/LRMs reveal a consistent gap between end-to-end image generation and reasoning-grounded construction. While UMMs excel at perceptual synthesis, they often fail to satisfy precise geometric constraints; in contrast, code-generating models demonstrate higher logical coherence and geometric correctness but remain limited in visual expressivity. We hope GGBench will serve as a rigorous foundation for future research in grounded, verifiable multimodal intelligence. Beyond geometry, our framework points toward a broader paradigm for evaluating generative reasoning—linking language, logic, and construction—as a step toward AI systems capable of building as well as understanding.